Find f(x) and g(x) so that the function can be described as y = f(g(x)).


y = 8/x^2 + 2



1
Expert's answer
2020-07-27T17:04:13-0400

Y=8/X2+2,

Y=2(4/X2+1),

Y=f(g(X)),

g(x)=4/x2+1,

f(x)=2x.